# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

### v3

@@ 3.1 Multi-phase gas kinematics
Figure 3 shows the moment maps of [Ne V], [Ne II], H2 S(1), and CO(3-2), obtained from the emission line fitting routines presented in Appendices C.1 and C.3. The [Ne V] seems to follow the direction of the ionized outflow, that is, the SE-NW direction (Venturi et al. 2018), as expected from species with high-IP (see Table LABEL:tab:_table_list_emission_lines). Its velocity field is also consistent with the [O III] kinematics, which is blueshifted in the SE and redshifted in the NW. In the SE (NW) cone, we estimate gas projected velocities up to -120 km s-1 (+100 km s-1). The velocity dispersion map of [Ne V] shows a donut-shaped structure in SE direction of the nucleus (highlighted with contours in Fig. 4), at a projected distance of 2.5″ ( 240 pc), with a peak of 170 km s-1.
The lower-IP [Ne II] emission line has a different morphology with respect to the highly ionized gas traced by the [Ne V], suggesting that it traces a different gas component. Indeed, the [Ne II] morphology resembles the stellar continuum emission in the circumnuclear ring (whose ALMA contours are highlighted in Fig. 1; see also Liu et al., 2023), as we can see in Fig. 3. Moreover, the moment-0 map reveals bright clumps and lanes in the left part of the FoV that match remarkably well the features visible in H map of Venturi et al. (2018). This supports the interpretation that the [Ne II] line predominantly traces the circumnuclear ring. This includes the elongated structure 4″ eastward of the nucleus, which is part of the so-called ‘Mid-East’ region - associated with the Southern Arm, a stream of gas flowing in from the southeast along the bar - and is interpreted as material moving downstream within the circumnuclear ring (Liu et al., 2023). The fact that [Ne II] traces the disk component is also confirmed by the velocity gradient of [Ne II], whose direction, along the galaxy disk major axis, and magnitude ( km s-1) are consistent with those of the rotating stellar component (and H) observed on the same scales by Venturi et al. (2018).
@@
The CO(3-2) moment maps shown in Fig. 3 are in agreement with the H2 S(1) kinematics, that is tracing the galaxy rotating disk. A similar CO velocity field was reported by Liu et al. (2023). This finding, which is consistent with the lack of any molecular outflow in the circumnuclear region (see also Combes et al., 2019), is in contrast to the CO (1-0) wind reported by Gao et al. (2021). Their outflow interpretation was based on velocity residuals from a rotating disk model on larger spatial scales, which resemble those observed in H and are more likely due to non-circular inflowing motions along the bar (see their Fig. 6). Notably, the cold gas traced by the CO(3-2) does not show the -enhanced region south of the nucleus as the H2, although both emission lines map the molecular gas. This suggests that the elevated velocity dispersion in H2 may trace non-circular motions, possibly related to local turbulence or streaming motions induced by the AGN, rather than a genuine outflow. In such an environment, the cold molecular gas traced by CO may not survive due to stronger dissociation or heating, leaving only the warm phase observable in this disturbed region.
The moment maps of the other emission lines listed in Table LABEL:tab:_table_list_emission_lines are shown in the online material, divided into molecular gas, recombination lines, and species with low ( 25 eV), medium ( 25 and 54 eV), and high ( 54 eV) IP. The thresholds are based on the ionization potentials of He I (25 eV) and He II (54 eV). Note that while some lines appear to lack emission in the nuclear region of the galaxy in the moment maps, this is due to low S/N where the spectra are dominated by strong continuum emission, rather than a genuine absence of emission. Grouping such emission lines by IP or gas phase leads to interesting typical patterns. All the H2 lines share kinematic features similar to those of H2 S(1) in Fig. 3, with a rotational disk kinematics and a high- blob to the South of the nucleus. The low-IP lines, as well as the recombination lines, show a velocity gradient consistent with a rotating disk, with an amplitude of 110 km s-1, like [Ne II]. The high- region observed in H2 is not present in all the low-IP and recombination lines. From the flux maps, we observed that these species are located mainly in the ‘Mid-East’ region (Liu et al., 2023). The moment maps of the high-IP species are similar to those of [Ne V] (see Fig. 3). In particular, the donut-shaped structure is well defined in the velocity dispersion maps, although in the Ch4 the spatial resolution is lower and the shape is not well recognizable. The medium-IP lines have intermediate properties compared to the previous two groups: the velocity gradient is similar to that of the low-IP lines, but the velocity dispersion maps show a high-velocity dispersion region in correspondence of the donut shape.
@@ 3.2 Velocity channel maps
Fig. 4 presents the velocity channel maps of the [O III] and [Ne V] emission lines, approximately spanning the velocity range from km s-1 to +550 km s-1 around the same systemic velocity. Both ionic species share the same kinematic structures in each velocity bin, although [Ne V] appears to trace regions closer to the outflow axis, as also found by Mingozzi et al. (2019) for high-ionization gas in MAGNUM galaxies. This is consistent with the fact that both trace the outflow but are associated with different IPs (97 eV and 35 eV for [Ne V] and [O III], respectively; see Table LABEL:tab:_table_list_emission_lines). Indeed, [Ne V] emission arises from more internal regions of the outflow, where the AGN radiation field is expected to be stronger. Interestingly, Fig. 3 shows that at projected velocities larger than 100 km s-1, the [Ne V] emission resembles the donut-shaped structure mentioned above. The bulk of the emission progressively shifts away from the nucleus with increasing channel velocity, tracing the lower arm of the donut-shaped structure. Moreover, unlike [O III], the redshifted [Ne V] velocity channels clearly reveal the receding NW ionization cone, which is otherwise highly obscured in the optical due to dust in the galaxy disk. Indeed, as discussed in the following, the NW cone lies behind the disk and suffers stronger extinction than the SE cone (Venturi et al., 2017, 2018; Marconcini et al., 2025b). Thanks to the lower dust attenuation in the mid-IR, the [Ne V] line is an optimal tracer of this obscured outflowing component.
@@ 3.3 Resolved diagnostic diagrams for gas excitation
Diagnostic diagrams featuring mid-IR emission line ratios are widely used to investigate the nature of the gas ionizing sources (e.g. Hao et al., 2009; Weaver et al., 2010; Inami et al., 2013; Richardson et al., 2022; Martínez-Paredes et al., 2023; Feltre et al., 2023; Garofali et al., 2024; Mingozzi et al., 2025). In particular, to investigate the gas excitation in the circumnuclear region of NGC 1365, we employed mid-IR line ratios recently proposed by Feltre et al. (2023), who analyzed integrated nuclear spectra of 42 local Seyfert galaxies observed with Spitzer, including NGC 1365. To ensure consistent comparisons between line ratios, we corrected the flux maps of each emission line for differences in FoV, pixel scale, and point spread functions (PSF). In practice, for each diagnostic diagram we adopted the FoV, pixel size, and PSF corresponding to the worst case among the lines involved, as detailed below. Additionally, as detailed in Appendix B, we applied multiplicative correction factors to account for flux discontinuities between the datacubes (see Fig. 11).
In the upper left panel of Fig. 5, we show the contours of line ratio model predictions for AGN (Feltre et al., 2023), SF galaxies (Gutkin et al., 2016), and shocks (Alarie & Morisset, 2019) for the [O IV]m/[Ne II]m and [Ne III]m/[Ne II]m line ratios. For this diagnostic diagram, we adopted the smallest FoV (6.6″ 7.7″, i.e. Ch3), the largest pixel size (0.35″ from Ch4), and convolved all images to the worst PSF (i.e. 1.0″, the one at the reddest wavelength of m). We include all available model grids, applying only a metallicity cut of (see Feltre et al. 2023 for further details on the models). We overlay our data on model predictions, where each pixel is color-coded based on its position in the corresponding diagnostic diagram. This diagram shows that the main ionization source in the circumnuclear region of NGC 1365 can either be shocks or AGN, while no evidence for SF excitation is detected. According to Fig. 5 of Feltre et al. (2023), adding fractional contributions from SF and shock models to an AGN model will shift the points towards the bottom-left in the diagram. Therefore, we classify as shock- or AGN-excited only those spaxels that fall entirely within a single predicted region. Data points located in overlapping or unclassified regions cannot be uniquely attributed to a single excitation mechanism. To better visualize the spatial distribution and morphology of these regions, we produce diagnostic maps using the same color-coding of the diagrams, as shown in the lower panels of Fig. 5.
@@
However, Laor (1998) showed that photoionizing shocks are extremely inefficient in powering the narrow line emission in AGN. Furthermore, as noted by Feltre et al. (2023), the predictions from shock models largely overlap with both AGN and star-formation grids in diagnostic diagrams. For these reasons, in Section 3.5 we consider only AGN and SF grids to model the optical and mid-IR emission lines.
Exploiting optical MUSE data, we compare the mid-IR diagnostic diagrams with the optical [S II] BPT diagram (Kewley et al., 2006, 2013) shown in the right column of Fig. 5. Moreover, we also considered the [N II] and [O I] BPT diagrams, finding similar results (see Fig. 5 in Venturi et al., 2018 and Appendix D3 in Mingozzi et al., 2019 for these diagrams at large scale). As found by Venturi et al. (2018), we notice that the circumnuclear SF ring seen in H and in CO (see Fig. 1) is not fully dominated by SF in the BPT diagrams. Instead, as visible from the [S II] spatial distribution, certain regions in the SE of the nucleus show signatures of AGN excitation. Venturi et al. (2018) conclude that this is likely due to the superposition along the line of sight of the AGN-ionized cone and the SF ring. The main difference between the optical BPT and the mid-IR diagnostic diagrams lies in their sensitivity to trace the true ionization source in dusty environments. In the nuclear region, the optical BPT appears to be dominated by star formation due to strong dust obscuration, as indicated by the Balmer decrement (see Appendix D and Fig. 12). In contrast, mid-IR diagnostics can penetrate the dust screen and clearly reveal AGN ionization at the nucleus. This explains why, in the optical, AGN ionization cones are only visible on larger spatial scales, as previously shown in Venturi et al. (2018) and Mingozzi et al. (2019). Another notable difference is that the region classified as shock-dominated in the mid-IR diagrams appears to be excited by the AGN radiation according to the optical diagnostic diagram. This discrepancy likely reflects the different depths probed by optical and mid-IR emission, with mid-IR lines tracing higher ionization regions deeper into the disk, where shocks from the interaction between the outflow and the disk are expected to occur.
@@ 3.4 Electron density estimation
One of the main advantages of MIRI is the availability of mid-IR density-sensitive line ratios, such as [Ne V]/[Ne V]. These ratios are sensitive to a higher density regime with respect to the commonly used optical [S II]/[S II] ratio, due to the different ionization energies and critical densities of the lines involved. Recent studies used IR tracers to explore the gas density in local Seyfert galaxies and found densities in the range log(ne/cm-3) = 3–5 (Hermosa Muñoz et al., 2024a, b; Zhang et al., 2024; Ceci et al., 2025; Ramos Almeida et al., 2025).
In the top panel of Fig. 6, we show the electron density map derived from the [Ne V]/[Ne V] ratio using PyNeb (Luridiana et al., 2015), assuming an electron temperature K (following typical values adopted for ionized gas in AGN narrow-line regions; Osterbrock & Ferland, 2006). Across the FoV, we estimate a median electron density of log(/cm-3) = (2.7 0.5), with a peak of log(/cm-3)3.5 at the nucleus. The density cannot be computed in some spaxels where the [Ne V] line ratio exceeds the lower density limit (LDL) of 1.2. In these spaxels, we can assume an upper limit for the density of log(ne/cm-3)2. In the remaining regions, the morphology of the density map appears to follow the conical structure of the outflow, as traced by the [Ne V] emission line (see Fig. 3).
Additionally, as shown in Appendix D, we computed the electron density map from the [S II] doublet observed in the MUSE data (see Venturi et al., 2018 and Mingozzi et al., 2019 for the maps in the entire MUSE FoV). Both the circumnuclear ring and the outflow region are characterized by peaks in electron density, with values reaching log(/cm-3)2.6–2.8. This spatial distribution is consistent with the findings of Venturi et al. (2018) and Kakkad et al. (2018), who reported enhanced electron densities in similar structures. For a quantitative comparison between the optical and mid-IR tracers, we measured the electron density from integrated spectra extracted within a 1.5″ ( 140 pc) radius aperture centered on the outflow region (see Fig. 6), using both the mid-IR and optical tracers. From the mid-IR [Ne V] line ratio, we estimate an electron density of (750440) cm-3, while from the [S II] lines we find a consistent value of (35040) cm-3. These values are illustrated in the bottom panel of Fig. 6, where we plot the theoretical [S II] and [Ne V] line ratios as functions of electron density using PyNeb (Luridiana et al., 2015), assuming an electron temperature K.
@@
Our attempt to compute from the [Ar V]/[Ar V] ratio within the 1.5″-radius aperture interestingly yielded a ratio of 2.1, which is above the [Ar V] LDL value of 1.8 (corresponding to an upper limit on log(ne/cm-3) density similar to the [Ne V] one, 2). More importantly, even when integrating the [Ne V] flux across the entire MIRI FoV, we do not observe line ratios falling in the low-density limit (LDL), contrary to what was found by Dudik et al. (2007) using Spitzer/IRS data. This further supports the idea that their higher [Ne V]/[Ne V] ratios - and apparent detection of LDL conditions - may originate from observational limitations. In particular, the higher sensitivity of JWST/MIRI MRS compared to Spitzer/IRS may have allowed for a more effective detection of low surface brightness emission that could have remained undetected in previous studies. A direct comparison with the Dudik et al. (2007) measurements and observational setup is presented in Appendix E.

### v4

@@ 3.1 Multiphase gas kinematics
Figure 3 shows the moment maps of [Ne V], [Ne II], H2 S(1), and CO(3-2), obtained from the emission line fitting routines presented in Appendices C.1 and C.3. The [Ne V] seems to follow the direction of the ionized outflow; that is, the SE-NW direction (Venturi et al. 2018), as expected from species with high-IP (see Table LABEL:tab:_table_list_emission_lines). Its velocity field is also consistent with the [O III] kinematics, which is blueshifted in the SE and redshifted in the NW. In the SE (NW) cone, we estimate gas projected velocities up to -120 km s-1 (+100 km s-1). The velocity dispersion map of [Ne V] shows a donut-shaped structure in SE direction of the nucleus (highlighted with contours in Fig. 4), at a projected distance of 2.5″ ( 240 pc), with a peak of 170 km s-1.
The lower-IP [Ne II] emission line has a different morphology with respect to the highly ionized gas traced by the [Ne V], suggesting that it traces a different gas component. Indeed, the [Ne II] morphology resembles the stellar continuum emission in the circumnuclear ring (whose ALMA contours are highlighted in Fig. 1; see also Liu et al., 2023), as we can see in Fig. 3. Moreover, the moment-0 map reveals bright clumps and lanes in the left part of the FoV that match remarkably well the features visible in H map of Venturi et al. (2018). This supports the interpretation that the [Ne II] line predominantly traces the circumnuclear ring. This includes the elongated structure 4″ eastward of the nucleus, which is part of the so-called “mid-east” region – associated with the Southern Arm, a stream of gas flowing in from the southeast along the bar – and is interpreted as material moving downstream within the circumnuclear ring (Liu et al., 2023). The fact that [Ne II] traces the disk component is also confirmed by the velocity gradient of [Ne II], whose direction, along the galaxy disk major axis, and magnitude ( km s-1) are consistent with those of the rotating stellar component (and H) observed on the same scales by Venturi et al. (2018).
@@
The CO(3-2) moment maps shown in Fig. 3 are in agreement with the H2 S(1) kinematics, which is tracing the galaxy rotating disk. A similar CO velocity field was reported by Liu et al. (2023). This finding, which is consistent with the lack of any molecular outflow in the circumnuclear region (see also Combes et al., 2019), is in contrast to the CO (1-0) wind reported by Gao et al. (2021). Their outflow interpretation was based on velocity residuals from a rotating disk model on larger spatial scales, which resemble those observed in H and are more likely due to noncircular inflowing motions along the bar (see their Fig. 6). Notably, the cold gas traced by the CO(3-2) does not show the -enhanced region south of the nucleus as the H2, although both emission lines map the molecular gas. This suggests that the elevated velocity dispersion in H2 may trace noncircular motions, possibly related to local turbulence or streaming motions induced by the AGN, rather than a genuine outflow. In such an environment, the cold molecular gas traced by CO may not survive due to stronger dissociation or heating, leaving only the warm phase observable in this disturbed region.
The moment maps of the other emission lines listed in Table LABEL:tab:_table_list_emission_lines are shown in the online material, divided into molecular gas, recombination lines, and species with low ( 25 eV), medium ( 25 and 54 eV), and high ( 54 eV) IP. The thresholds are based on the IPs of He I (25 eV) and He II (54 eV). Note that while some lines appear to lack emission in the nuclear region of the galaxy in the moment maps, this is because the spectra are dominated by strong continuum emission rather than a genuine absence of emission. Grouping such emission lines by IP or gas phase leads to interesting typical patterns. All the H2 lines share kinematic features similar to those of H2 S(1) in Fig. 3, with a rotational disk kinematics and a high- blob to the south of the nucleus. The low-IP lines, as well as the recombination lines, show a velocity gradient consistent with a rotating disk, with an amplitude of 110 km s-1, like [Ne II]. The high- region observed in H2 is not present in all the low-IP and recombination lines. From the flux maps, we observed that these species are located mainly in the mid-east region (Liu et al., 2023). The moment maps of the high-IP species are similar to those of [Ne V] (see Fig. 3). In particular, the donut-shaped structure is well defined in the velocity dispersion maps, although in the Ch4 the spatial resolution is lower and the shape is not well recognizable. The medium-IP lines have intermediate properties compared to the previous two groups: the velocity gradient is similar to that of the low-IP lines, but the velocity dispersion maps show a high-velocity dispersion region in correspondence of the donut shape.
@@ 3.2 Velocity channel maps
Fig. 4 presents the velocity channel maps of the [O III] and [Ne V] emission lines, approximately spanning the velocity range from 450 km s-1 to +550 km s-1 around the same systemic velocity. Both ionic species share the same kinematic structures in each velocity bin, although [Ne V] appears to trace regions closer to the outflow axis, as also found by Mingozzi et al. (2019) for high-ionization gas in MAGNUM galaxies. This is consistent with the fact that both trace the outflow but are associated with different IPs (97 eV and 35 eV for [Ne V] and [O III], respectively; see Table LABEL:tab:_table_list_emission_lines). Indeed, [Ne V] emission arises from more internal regions of the outflow, where the AGN radiation field is expected to be stronger. Interestingly, Fig. 3 shows that at projected velocities larger than 100 km s-1, the [Ne V] emission resembles the donut-shaped structure mentioned above. The bulk of the emission progressively shifts away from the nucleus with increasing channel velocity, tracing the lower arm of the donut-shaped structure. Moreover, unlike [O III], the redshifted [Ne V] velocity channels clearly reveal the receding NW ionization cone, which is otherwise highly obscured in the optical due to dust in the galaxy disk. Indeed, as is discussed in the following, the NW cone lies behind the disk and suffers stronger extinction than the SE cone (Venturi et al., 2017, 2018; Marconcini et al., 2025b). Thanks to the lower dust attenuation in the mid-IR, the [Ne V] line is an optimal tracer of this obscured outflowing component.
@@ 3.3 Resolved diagnostic diagrams for gas excitation
Diagnostic diagrams featuring mid-IR emission line ratios are widely used to investigate the nature of the gas ionizing sources (e.g., Hao et al., 2009; Weaver et al., 2010; Inami et al., 2013; Richardson et al., 2022; Martínez-Paredes et al., 2023; Feltre et al., 2023; Garofali et al., 2024; Mingozzi et al., 2025). In particular, to investigate the gas excitation in the circumnuclear region of NGC 1365, we employed mid-IR line ratios recently proposed by Feltre et al. (2023), who analyzed integrated nuclear spectra of 42 local Seyfert galaxies observed with Spitzer, including NGC 1365. To ensure consistent comparisons between line ratios, we corrected the flux maps of each emission line for differences in FoV, pixel scale, and point spread functions (PSFs). In practice, for each diagnostic diagram we adopted the FoV, pixel size, and PSF corresponding to the worst case among the lines involved, as is described below. Additionally, as is detailed in Appendix B, we applied multiplicative correction factors to account for flux discontinuities between the datacubes (see Fig. 11).
In the upper left panel of Fig. 5, we show the contours of line ratio model predictions for AGN (Feltre et al., 2023), SF galaxies (Gutkin et al., 2016), and shocks (Alarie and Morisset, 2019) for the [O IV]m/[Ne II]m and [Ne III]m/[Ne II]m line ratios. For this diagnostic diagram, we adopted the smallest FoV (6.6″ 7.7″, i.e., Ch3), the largest pixel size (0.35″ from Ch4), and convolved all images to the worst PSF (i.e., 1.0″, the one at the reddest wavelength of m). We include all available model grids, applying only a metallicity cut of (see Feltre et al. 2023 for further details on the models). We overlay our data on model predictions, where each pixel is color-coded based on its position in the corresponding diagnostic diagram. This diagram shows that the main ionization source in the circumnuclear region of NGC 1365 can either be shocks or AGN, while no evidence for SF excitation is detected. According to Fig. 5 of Feltre et al. (2023), adding fractional contributions from SF and shock models to an AGN model will shift the points toward the bottom left in the diagram. Therefore, we classify as shock- or AGN-excited only those spaxels that fall entirely within a single predicted region. Data points located in overlapping or unclassified regions cannot be uniquely attributed to a single excitation mechanism. To better visualize the spatial distribution and morphology of these regions, we produced diagnostic maps using the same color-coding of the diagrams, as is shown in the lower panels of Fig. 5.
@@
However, Laor (1998) showed that photoionizing shocks are extremely inefficient in powering the narrow line emission in AGNs. Furthermore, as has been noted by Feltre et al. (2023), the predictions from shock models largely overlap with both AGN and star-formation grids in diagnostic diagrams. For these reasons, in Section 3.5 we consider only AGN and SF grids to model the optical and mid-IR emission lines.
Exploiting optical MUSE data, we compare the mid-IR diagnostic diagrams with the optical [S II] BPT diagram (Kewley et al., 2006, 2013) shown in the right column of Fig. 5. Moreover, we also considered the [N II] and [O I] BPT diagrams, finding similar results (see Fig. 5 in Venturi et al., 2018 and Appendix D3 in Mingozzi et al., 2019 for these diagrams at large scale). As found by Venturi et al. (2018), we notice that the circumnuclear SF ring seen in H and in CO (see Fig. 1) is not fully dominated by SF in the BPT diagrams. Instead, as visible from the [S II] spatial distribution, certain regions in the SE of the nucleus show signatures of AGN excitation. Venturi et al. (2018) conclude that this is likely due to the superposition along the line of sight of the AGN-ionized cone and the SF ring. The main difference between the optical BPT and the mid-IR diagnostic diagrams lies in their sensitivity to trace the true ionization source in dusty environments. In the nuclear region, the optical BPT appears to be dominated by SF due to strong dust obscuration, as is indicated by the Balmer decrement (see Appendix D and Fig. 12). In contrast, mid-IR diagnostics can penetrate the dust screen and clearly reveal AGN ionization at the nucleus. This explains why, in the optical, AGN ionization cones are only visible on larger spatial scales, as previously shown in Venturi et al. (2018) and Mingozzi et al. (2019). Another notable difference is that the region classified as shock-dominated in the mid-IR diagrams appears to be excited by the AGN radiation according to the optical diagnostic diagram. This discrepancy likely reflects the different depths probed by optical and mid-IR emission, with mid-IR lines tracing higher ionization regions deeper into the disk, where shocks from the interaction between the outflow and the disk are expected to occur.
@@ 3.4 Electron density estimation
One of the main advantages of MIRI is the availability of mid-IR density-sensitive line ratios, such as [Ne V]/[Ne V]. These ratios are sensitive to a higher density regime with respect to the commonly used optical [S II]/[S II] ratio, due to the different ionization energies and critical densities of the lines involved. Recent studies used IR tracers to explore the gas density in local Seyfert galaxies and found densities in the range log(ne/cm-3) = 3–5 (Hermosa Muñoz et al., 2025, 2024; Zhang et al., 2024; Ceci et al., 2025; Ramos Almeida et al., 2025).
In the top panel of Fig. 6, we show the electron density map derived from the [Ne V]/[Ne V] ratio using PyNeb (Luridiana et al., 2015), assuming an electron temperature K (following typical values adopted for ionized gas in AGN narrow-line regions; Osterbrock and Ferland, 2006). Across the FoV, we estimate a median electron density of log(/cm-3) = (2.7 0.5), with a peak of log(/cm-3)3.5 at the nucleus. The density cannot be computed in some spaxels where the [Ne V] line ratio exceeds the lower density limit (LDL) of 1.2. In these spaxels, we can assume an upper limit for the density of log(ne/cm-3)2. In the remaining regions, the morphology of the density map appears to follow the conical structure of the outflow, as traced by the [Ne V] emission line (see Fig. 3).
Additionally, as is shown in Appendix D, we computed the electron density map from the [S II] doublet observed in the MUSE data (see Venturi et al., 2018 and Mingozzi et al., 2019 for the maps in the entire MUSE FoV). Both the circumnuclear ring and the outflow region are characterized by peaks in electron density, with values reaching log(/cm-3)2.6–2.8. This spatial distribution is consistent with the findings of Venturi et al. (2018) and Kakkad et al. (2018), who reported enhanced electron densities in similar structures. For a quantitative comparison between the optical and mid-IR tracers, we measured the electron density from integrated spectra extracted within a 1.5″ ( 140 pc) radius aperture centered on the outflow region (see Fig. 6), using both the mid-IR and optical tracers. From the mid-IR [Ne V] line ratio, we estimate an electron density of (750440) cm-3, while from the [S II] lines we find a consistent value of (35040) cm-3. These values are illustrated in the bottom panel of Fig. 6, where we plot the theoretical [S II] and [Ne V] line ratios as functions of electron density using PyNeb (Luridiana et al., 2015), assuming an electron temperature K.
@@
Our attempt to compute from the [Ar V]/[Ar V] ratio within the 1.5″-radius aperture interestingly yielded a ratio of 2.1, which is above the [Ar V] LDL value of 1.8 (corresponding to an upper limit on log(ne/cm-3) density similar to the [Ne V] one, 2). More importantly, even when integrating the [Ne V] flux across the entire MIRI FoV, we do not observe line ratios falling in the LDL, contrary to what was found by Dudik et al. (2007) using Spitzer/IRS data. This further supports the idea that their higher [Ne V]/[Ne V] ratios - and apparent detection of LDL conditions - may originate from observational limitations. In particular, the higher sensitivity of JWST/MIRI MRS compared to Spitzer/IRS may have allowed for a more effective detection of low surface brightness emission that could have remained undetected in previous studies. A direct comparison with the Dudik et al. (2007) measurements and observational setup is presented in Appendix E.